with the following chemical reaction, if we start with 6.81 grams of magnesium, how many litres of oxygen are required? 2Mg + O2 = 2MgO
1
Expert's answer
2020-05-28T10:32:17-0400

2Mg + O2 = 2MgO

n(Mg) = 6.81/24 = 0.2838 mol

n(O2) = n(Mg)/2 = 0.2838/2 = 0.1419 mol

V(O2) = 0.1419*22.4 = 3.18 L